[Date Prev][Date Next][Thread Prev][Thread Next] - [Date Index][Thread Index][Author Index]

Opal Update - JAK and STENSAT have left the nest

Opal Update: 2/11/2000

- Last night, fire commands were sent to Opal to launch
  tube 4 containing JAK and STENSAT.   Opal telemetry
  did not confirm or deny the launch.  Difficulities in
  the connection resulted in corrupted data downloads.

  During the AM pass today, we established a connection
  and downloaded pico launch tube telemetry.  Data
  indicates SUCCESSFULL PICO LAUNCH!  Two data sets
  confirm this.   Sorry for the long URLs.

        ROS 11:

        Solar panel:

  ROS 11 is a reflective object sensor in tube 4 that
  detects the pico.  A high voltage indicates absence
  of picos, a low indicates presence of picos.  You'll
  notice the plot starts high and they dips down.
  This is a "feature" that we fully expected.  Light from
  the sun produces a reflection signal thus tricking our
  sensor.  The solar panel plot confirms that the pico
  doors are in the sun at the time of the decreased signal.
  The doors are on side 2.  Solar panel 7 and 8 are top
  and bottom respectively.  1 and 3 are on either side of
  the doors.

  We are confident this data indicates pico launch.  We hve
  not checked with sensors in Tube 1 and 2 that still house
  picos. If we have time, we will.

  Space Command is monitoring Opal and looking for picos.
  We should have the data within several hours.

  So, JAK and STENSAT are space-borne!   Go get em guys.

- We are complete pleased at the contacts these past two
  days.  However, we are concerned about battery temperature
  data.  Our recent data indicates temperatures in the range
  50-70C.  This far exceeds their ratings and could result
  in premature failure.  We believe this is due to over-
  charging of the batteries.  We have turned on several Opal
  payload systems to increase our consumption of power.
  On the next pass, we will download more spacecraft telemetry
  to that will aid us in characterizing Opal.  Our recent
  data sets are only several point measurements, rather
  than continuous data measurements.  We hope to download some
  10 hour data sets tonight.

- The spurious tone is still periodically present on Opal's
  transmissions.  However, the transmissions are appearing
  at hte appropriate times and lengths.  This week we entered
  a mode where the AX 25 beacon reduced its frequency from
  6 times a minute to once a minute.  This was due to a
  decrease in battery voltage.  The CPU no longer was alarmed
  and therefore decreased the frequency of beacons.  We lowered
  the threshold alarm voltage to enable the increased beacon
  frequency.  This will help us in tracking and also in
  dissipating power.  Unfortunately this could mask the SCU
  picos that are on the same frequency.  We will attempt to
  lower the frequency again when we launch the last two.

  What's interesting is that this morning we were able to
  connect and talk to Opal despite the presence of the spurious
  tone.  It seemed to be higher pitch than before.  In fact,
  the signals from Opal sounded quite bad but we decoded them
  perfectly.  Gotta like that big dish.  Howevery, other
  people were also able to decode packets this AM.

- Two more SCU picos are still housed in Opal.  These are the
  VLF lightning picos.  We expect to launch them this PM or
  tomorrow AM.

- Our time at the dish ends tomorrow evening.  We will return
  our equipment to Stanford an attempt to establish contact
  from here.  Several HAMs have been decoding our recent
  contact sessions from the dish and have heard Opal clearly.
  This  gives us hope that our station will indeed be able
  to comm with Opal.  It's goign to take a lot of work though.
  Too bad we can't use the dish on a continued basis.

- Thanks for your continued support and encouragement.  I think
  we're all amazed that the mission is progressing so well!

Via the amsat-bb mailing list at AMSAT.ORG courtesy of AMSAT-NA.
To unsubscribe, send "unsubscribe amsat-bb" to Majordomo@amsat.org